OpenStudy (anonymous):

how do I differentiate 5/x^4

OpenStudy (anonymous):

Write it as 5 * x^-4. Differentiate using power rule, and presto.

myininaya (myininaya):

5/(x^4)=5x^(-4) (5x^(-4))'=-4*5x^(-4-1)